An object is moving on a circular track of radius r with constant speed V. The avg acc of object during time magnitude of interval  in  which it complete One by fourth revolution of the track will be.... 

Grade:12th pass

1 Answers

Khimraj
5 years ago
centripital acceleration ac = V2/r

Tangential accl.(at) is equal to

Change in velocity/ time taken.

change in velocity = V\sqrt{2}

time taken = (2\pi r/4)/V = \pi r/2V

at = 2\sqrt{2}V^{2}/\pi r

average acceleration a = \sqrt{a_{c}^{2}+ a_{t}^{2}}